OpenStudy (anonymous):

What percent of 51 is 127.5?

TheSmartOne (thesmartone):

So that means (x/100) x 51=127.5

OpenStudy (abb0t):

Why don't you try this one yourself instead of asking the same question over, and over, and over again.

OpenStudy (anonymous):

Why don't you go kill yourself? This is my account and I can do whatever the f*uck I want

TheSmartOne (thesmartone):

\[\frac{ x }{ 100 } \times 51 =127.5\]

OpenStudy (abb0t):

Doing it yourself and messing up is part of the learning process, kid. So try it yourself!

TheSmartOne (thesmartone):

Ok now I am not going to help you. We are all here to help you. We want you to learn it and then remember it so that it can help you for the rest of your life.

TheSmartOne (thesmartone):

And there is no need to use profanity...

OpenStudy (tkhunny):

Yup. Try it yourself when your brief suspension is over. How about that? I agreed with bibby one day and abb0t the next. :-)